# Generalised rank-constrained approximations of Hilbert–Schmidt operators on separable Hilbert spaces and applications

### v1

@@ Abstract
$\min\{\lVert M-BXC\rVert_{HS}:\ \dim\operatorname{ran}{X}\leq r\}.$
This extends the result of Sondermann (Statistische Hefte, 1986) and Friedland and Torokhti (SIAM J. Matrix Analysis and Applications, 2007), which studies this problem in the case of matrices $M$, $B$, $C$, $X$, and the analysis involves the Moore–Penrose inverse. In classical approximation problems which can be solved by the singular value decomposition or Moore–Penrose inverse, the solution satisfies a minimal norm property. Friedland and Torokhti also state such a minimal norm property of the solution. However, we show that this minimal norm property does not hold in general and give a modified minimality property that is satisfied in general. We also show that the solution may be discontinuous in infinite-dimensional settings. Necessary and sufficient conditions for continuity of the solutions are given and continuous approximations are constructed when such conditions are not met. Finally, we show that our results can be used in signal processing, reduced rank regression and linear operator learning under a rank constraint.
@@ 1 Introduction
We study the following reduced rank approximation in Hilbert–Schmidt norm $\lVert\cdot\rVert_{HS}$,
$\min\{\lVert M-BXC\rVert_{HS}:\ \dim\operatorname{ran}{X}\leq r\},$ (1)
@@
In the finite-dimensional matrix case, the generalised approximation problem is well studied, cf. [38, 20, 40]. In [20, Theorem 2.1] and [38] a solution of the finite-dimensional matrix problem is given. Furthermore, conditions for uniqueness of the solution are given. A minimality property of the solution is also given in [20, Theorem 2.1], namely that it should have minimal norm among all solutions of the approximation problem. In [40] an error analysis of the solution proposed by [20, Theorem 2.1] is given.
@@ 1.1 Contributions of this work
The first aim of this work is to show that a conclusion in [20, Theorem 2.1] — namely, that the solution of a certain generalised approximation problem satisfies a minimal norm property — is not in general correct. We therefore propose a refinement of [20, Theorem 2.1] by modifying the minimal norm property. This modified property is always satisfied by the solution given in [20, Theorem 2.1] and can still be interpreted as a minimality property.
The second aim of this work is to generalise the refined statement to an infinite-dimensional setting. We show that the generalised approximation problem does on the one hand not always admit an optimal solution. On the other hand, even when solutions exist, they are no longer bounded in general, due to the Moore–Penrose inverse of $C$ that appears in the solution expression. It is even possible that some solutions are continuous, i.e. bounded, and have finite Hilbert–Schmidt norm, while others are discontinuous, i.e. unbounded, and as such do not even have finite operator norm. Formulation and analysis of the problem thus requires the use of unbounded operators. While the Moore–Penrose inverse is defined also for certain kinds of unbounded operators such as closed range operators, we analyse the problem assuming that $B$ and $C$ do not only have closed range but are bounded linear operators. Two natural questions arise: when does a solution exist, and when are the solutions bounded? We give necessary and sufficient conditions for a solution to the generalised approximation problem to exist in infinite dimensions. When it does not exist, i.e. when the infimum is not attained, we construct a sequence of approximate solutions that converges to this infimum. We shall also give necessary and sufficient conditions for all solutions to be bounded. When no bounded solutions exist, we propose a procedure to construct bounded approximations.
@@
The third aim of this work is to demonstrate how our results can be used in relevant applications of current research interest such as operator learning [30] and in particular linear operator learning [17, 36], reduced rank regression [27, 28, 62] and signal processing [51, 54]. In such applications, there is a direct interest in an infinite-dimensional formulation. For example, as mentioned in [36], the setting of infinite-dimensional linear operator learning has direct applications to functional regression with functional response and conditional mean embeddings. In such an operator learning setting, given samples from the distribution of Hilbert space-valued random variables $x$ and $y$, the goal is to learn a linear operator $A$ that satisfies
$\min\{\mathbb{E}\lVert y-Ax\rVert^{2},A\in\mathcal{A}\},$ (2)
@@
where $\mathcal{A}$ can equal for example the bounded linear or Hilbert–Schmidt operators. This is also studied as a linear signal processing problem in [51], where $x$ and $y$ are $\mathbb{R}^{m}$ or $\mathbb{C}^{m}$ valued. In [51] $x$, $y$, and $A$ are also allowed to be reweighted and the rank of $A$ can be restricted. Under such a rank restriction, the above problem can in fact also be seen as a reduced rank regression problem similar to [27, 62]. We show that these problems can be rewritten in formulation (1). In [28], a reduced rank regression problem of the form (1) is considered where the Hilbert–Schmidt norm is replaced by the operator norm. Applications from the perspective of linear operator learning, signal processing and reduced rank regression thus all motivate the study of a problem of form (2). Using the obtained solution of problem (1) in the infinite-dimensional setting, we therefore analyse problem (2), where $\mathcal{A}$ is the space of finite-rank operators between separable Hilbert spaces. We also allow for reweighting of $x,y$ and $A$ as in [51], and give necessary and sufficient conditions for a solution to exist, constructing also an explicit solution that satisfies a minimality property. In this application, we show that our proposed minimality property of the solutions corresponds to a maximal kernel property that naturally occurs in linear operator learning.
@@
Even in finite-dimensional settings, the study of an infinite-dimensional analogue can be beneficial. Infinite-dimensional problems are often discretized and transformed to finite-dimensional generalised rank-constrained approximation problems. This happens for example in the study of inverse problems governed by partial differential equations, e.g. [12]. An infinite-dimensional version of the generalised approximation problem would then enable the analysis of such problems directly in their infinite-dimensional formulation, enabling us to draw conclusions that are independent of the discretization dimension.
@@ 1.2 Related literature
Imposing a minimality property has the consequence of mitigating nonuniqueness of the solution. This can also be achieved by imposing a regularizing term $\alpha^{2}\lVert X\rVert_{HS}^{2}$. This is the approach of [13] for a similar reduced rank approximation problem in the context of inverse problems.
@@ 1.3 Notation
We let $\mathcal{H}$ and $\mathcal{K}$ be possibly infinite-dimensional separable Hilbert spaces over $\mathbb{C}$ or $\mathbb{R}$. By $\mathcal{B}(\mathcal{H}$, $\mathcal{K})$, $\mathcal{B}_{0}(\mathcal{H}$, $\mathcal{K})$, $\mathcal{B}_{00}(\mathcal{H}$, $\mathcal{K})$ and $L_{2}(\mathcal{H},\mathcal{K})$, we denote the linear spaces of bounded, compact, finite-rank and Hilbert–Schmidt operators between Hilbert spaces $\mathcal{H}$ and $\mathcal{K}$. The Hilbert–Schmidt norm of a linear operator between finite-dimensional spaces is equal to the Frobenius norm of any matrix representation of this operator in an orthonormal basis. By $\mathcal{B}_{00,r}(\mathcal{H},\mathcal{K})$ we denote set of finite-rank operators, i.e. operators that are bounded, have finite-dimensional range, and have rank at most $r\in\mathbb{N}$. Further details of these classes of operators is given in Appendix A.
@@ 2 Optimal approximation for matrices
$\min{\left\{\lVert M-BXC\rVert_{F},\ X\in\mathbb{C}^{q\times p},\ \operatorname{rank}\left({X}\right)\leq r\right\}},$ (4)
@@
A few remarks are in order. Firstly, in general not all solutions of (4) are of the form (3). Secondly, as $P_{C,R}=C^{\dagger}C$ and as in finite dimensions it holds that $P_{B,L}=BB^{\dagger}$, one may write (3) as $B^{\dagger}(BB^{\dagger}XC^{\dagger}C)_{r}C^{\dagger}$. Thirdly, also a necessary and sufficient condition for uniqueness of $\hat{X}$ is given in [20, Theorem 2.1]. Finally, if $(P_{B,L}MP_{C,R})_{r}$ is not uniquely specified, then the matrix $\hat{X}$ in (3) is not uniquely determined. Hence eq. 3 should be interpreted as a set of solutions. Furthermore, this lack of uniqueness extends also to $\lVert\hat{X}\rVert_{HS}$: if there exists more than one solution of the form (3), then these solutions might not all have the same norm. In other words, whether $\hat{X}$ in (3) is of minimal norm or not in general depends on the choice of truncated SVD $(P_{B,L}MP_{C,R})_{r}$. This fact indicates that the claim above is not sufficiently precise.
@@
So $M=P_{B,L}P_{C,R}$ and its the rank-1 truncated SVDs are,
@@
The two minimizing matrices of the form (3) then are,
$\hat{X}_{a}\coloneqq\begin{pmatrix}1&0&0\\ 0&0&0\\ 0&0&0\\ \end{pmatrix},\quad\hat{X}_{b}\coloneqq\begin{pmatrix}0&0&0\\ 0&4&0\\ 0&0&0\\ \end{pmatrix},$
@@
The preceding example suggests that [20, Theorem 2.1] can be made more precise, either by removing the minimum norm condition (5), or by replacing this condition by a weaker condition which all solutions $\hat{X}$ given by (3) satisfy. Such a weaker condition must thus be independent of the choice of truncated SVD $(P_{B,L}MP_{C,R})_{r}$. Furthermore, if $C$ is a linear operator mapping between infinite-dimensional spaces, then ${C}^{\dagger}$ will in general not be bounded (cf. Proposition A.5), and thus cannot be Hilbert–Schmidt. It can then happen that also solutions $\hat{X}$ of the form (3) are not Hilbert–Schmidt, in which case $\lVert\hat{X}\rVert_{HS}$ may be ill-defined, and we need a minimality condition that does not involve the Frobenius norm in order to generalise this condition to the infinite-dimensional setting. For these two reasons we motivate an alternative condition, by continuing our Example 2.1.
@@
Like $B\hat{X}_{c}(\alpha)C$, also $BX(\alpha)C$ does not depend on $\alpha$. By the Eckart–Young Theorem (c.f. Proposition A.7), for $X(\alpha)$ to be a solution of (4), we need $BX(\alpha)C$ to be a rank-1 SVD of $M$. That is, we need $x_{12}=x_{21}=0$ and either $x_{11}=1$ or $x_{22}=4$.
@@
Let $e_{i}$ denote the $i$-th canonical basis vector of $\mathbb{C}^{3}$ and $V\coloneqq\operatorname{span}{\left(e_{1},e_{2}\right)}\subset\mathbb{C}^{3}$. Let $P_{V}$ the matrix representation of the projector onto $V$ in the canonical basis of $\mathbb{C}^{3}$. Then for $B,C$ defined in (6), $P_{B,R}=P_{V}=P_{C,L}$. We observe $P_{B,L}\hat{X}(\alpha)P_{C,R}=P_{V}\hat{X}_{c}(\alpha)P_{V}=\hat{X}_{c}(0)$ for any $\alpha$, and compute,
$\lVert\hat{X}_{c}(\alpha)\rVert_{HS}^{2}=\lVert X_{c,11}\rVert_{HS}^{2}+\sum_{i=1}^{5}\lvert\alpha_{i}\rvert^{2}\geq\lVert X_{c,11}\rVert_{HS}^{2}=\lVert\hat{X}_{c}(0)\rVert_{HS}^{2}.$ (7)
@@
that are the minimisers in (7) for some $c\in\{a,b\}$. Hence (8) provides an alternative for the minimal norm property (5). So, while $X_{a}(0)$ and $X_{b}(0)$ might not have the same norm, $X_{c}(0)$ has minimal norm among all $X_{c}(\alpha),\alpha\in\mathbb{C}^{5}$, for fixed $c\in\{a,b\}$, and $X_{a}(0),X_{b}(0)$ are in turn those solutions of (4) which satisfy (8).
To generalise the reasoning of Example 2.1 as to why (8) can be interpreted as a minimality condition, notice that (8) is equivalent to $(I-P_{B,R})\hat{X}P_{C,L}=0$ and $\hat{X}(I-P_{C,L})=0$, since any $X\in\mathbb{C}^{q\times p}$ can be decomposed as
@@
If $\hat{X}$ in (3) is unique, then $\lVert\hat{X}\rVert_{HS}$ is indeed minimal among all minimisers $X$ of (4), in which case the statement of [20, Theorem 2.1] is still valid. Of course, this holds only when Hilbert–Schmidt norms can in fact be computed, which is no longer true in general for infinite-dimensional spaces. In the next section we introduce a statement similar to [20, Theorem 2.1] with the modified minimality property (8) in an infinite-dimensional setting.

### v3

@@ Abstract
$\min\{\lVert M-BXC\rVert_{L_{2}}:\ \dim\operatorname{ran}{X}\leq r\}.$
This extends the result of Sondermann (Statistische Hefte, 1986) and Friedland and Torokhti (SIAM J. Matrix Analysis and Applications, 2007), which studies this problem in the case of matrices $M$, $B$, $C$, $X$, and the analysis involves the Moore–Penrose inverse. In classical approximation problems that can be solved by the singular value decomposition or Moore–Penrose inverse, the solution satisfies a minimal norm property. Friedland and Torokhti state such a minimal norm property of the solution. We show that this minimal norm property does not hold in general and give a modified minimality property that does hold. We show that the solution may be discontinuous in infinite-dimensional settings. We give conditions for continuity of the solutions and construct continuous approximations when such conditions are not met. Finally, we study problems from signal processing, reduced rank regression and linear operator learning under a rank constraint. Our theoretical results enable us to explicitly find solutions to these problems and to characterise their existence, uniqueness and minimality property.
@@ 1 Introduction
We study the following reduced rank approximation in Hilbert–Schmidt norm $\lVert\cdot\rVert_{L_{2}}$,
$\min\{\lVert M-BXC\rVert_{L_{2}}:\ \dim\operatorname{ran}{X}\leq r\},$ (1)
@@
In the finite-dimensional matrix case, the generalised approximation problem is well studied, cf. [38, 20, 40]. In [20, Theorem 2.1] and [38] a solution of the finite-dimensional matrix problem is formulated. Furthermore, conditions for uniqueness of the solution are given. A minimality property of the solution is also stated in [20, Theorem 2.1], namely that it should have minimal norm among all solutions of the approximation problem. In [40] an error analysis of the solution proposed by [20, Theorem 2.1] is added.
@@ 1.1 Contributions of this work
Firstly, the theoretical contributions of this work are as follows:
  • We show in Section 2 that a conclusion in [20, Theorem 2.1]—namely, that the solution of (1) in a finite-dimensional matrix formulation satisfies a minimal norm property—is not in general correct. In many applications, the minimal norm property is often seen as a desirable property of the solution.
@@
  • We propose a refinement of [20, Theorem 2.1] by modifying the minimal norm property in Section 2. This modified property is always satisfied by the solution given in [20, Theorem 2.1] and can still be interpreted as a minimality property.
  • We generalise the refined statement to an infinite-dimensional setting Theorem 3.2, and note that the infinite-dimensional approximation problem does not always admit an optimal solution by giving a necessary and sufficient existence condition.
  • In the case where the generalised approximation problem (1) admits one or more optimal solutions, we characterise the solutions that satisfy the modified minimality property in Theorem 3.2, as well as the solutions that do not satisfy the modified minimality property in Corollary 3.9. In the case where (1) does not admit optimal solutions, we construct a sequence of approximate minimisers of the objective $\lVert M-BXC\rVert_{L_{2}}$ in Section 3.2.
  • We calculate the minimal approximation error that can be attained for the problem (1) in Section 3.3.
  • We show that solutions of (1) are not bounded in general in Section 4, due to the Moore–Penrose inverse of $C$ that appears in the solution expression. It is even possible that some solutions are continuous, i.e. bounded, and have finite Hilbert–Schmidt norm, while others are discontinuous, i.e. unbounded, and as such do not even have finite operator norm.
  • We give several necessary or sufficient conditions for all solutions of (1) to be bounded in Section 5.1.
  • In the case where the problem (1) does not admit any bounded solutions, we propose a procedure to construct bounded approximations in Section 5.2. This approximation procedure also allows us to formulate an adjoint problem to (1), which in certain cases can be useful to recover bounded approximations.
While the Moore–Penrose inverse is defined for certain kinds of unbounded operators such as closed range operators, we restrict our analysis to the setting where $B$ and $C$ not only have closed range, but are bounded linear operators.
@@
The second aim of this work is to demonstrate how our results can be used in relevant applications of current research interest such as operator learning [30] and in particular linear operator learning [17, 36], reduced rank regression [27, 28, 62] and signal processing [51, 54]. In such applications, there is a direct interest in an infinite-dimensional formulation. For example, as mentioned in [36], the setting of infinite-dimensional linear operator learning has direct applications to functional regression with functional response and conditional mean embeddings. In such an operator learning setting, given samples from the distribution of Hilbert space-valued random variables $x$ and $y$, the goal is to learn a linear operator $A$ that satisfies
$\min\{\mathbb{E}\lVert y-Ax\rVert^{2}:A\in\mathcal{A}\},$ (2)
@@
where $\mathcal{A}$ can be for example the bounded linear or Hilbert–Schmidt operators. This is also studied as a linear signal processing problem in [51], where $x$ and $y$ are $\mathbb{R}^{m}$ or $\mathbb{C}^{m}$ valued. In [51] $x$, $y$, and $A$ are also allowed to be reweighted and the rank of $A$ can be restricted. Under such a rank restriction, the above problem can in fact also be seen as a reduced rank regression problem similar to [27, 62]. We show that these problems can be rewritten in formulation (1). In [28], a reduced rank regression problem of the form (1) is considered where the Hilbert–Schmidt norm is replaced by the operator norm. Applications from the perspective of linear operator learning, signal processing and reduced rank regression thus all motivate the study of a problem of form (2). Using the obtained solution of problem (1) in its adjoint formulation in the infinite-dimensional setting, we therefore analyse problem (2), where $\mathcal{A}$ is the space of finite-rank operators between separable Hilbert spaces. We also allow for reweighting of $x,y$ and $A$ as in [51], and give necessary and sufficient conditions for a solution to exist, constructing also an explicit solution that satisfies a minimality property. In this application, we show that our proposed minimality property of the solutions corresponds to a maximal kernel property that naturally occurs in linear operator learning, c.f. [36, Remark 3.12(b)].
@@
Even in finite-dimensional settings, the study of an infinite-dimensional analogue can be beneficial. Infinite-dimensional problems are often discretised and transformed to finite-dimensional generalised rank-constrained approximation problems. This happens for example in the study of inverse problems governed by partial differential equations, e.g. [12]. An infinite-dimensional version of the generalised approximation problem would then enable the analysis of such problems directly in their infinite-dimensional formulation, allowing us to draw conclusions that are independent of the discretisation dimension.
@@ 1.2 Related literature
Imposing a minimality property has the consequence of mitigating nonuniqueness of the solution. This can also be achieved by imposing a regularizing term $\alpha^{2}\lVert X\rVert_{L_{2}}^{2}$. This is the approach of [13] for a similar reduced rank approximation problem in the context of inverse problems.
@@ 1.3 Notation
We let $\mathcal{H}$ and $\mathcal{K}$ be possibly infinite-dimensional separable Hilbert spaces over $\mathbb{C}$ or $\mathbb{R}$. By $\mathcal{B}(\mathcal{H}$, $\mathcal{K})$, $\mathcal{B}_{0}(\mathcal{H}$, $\mathcal{K})$, $\mathcal{B}_{00}(\mathcal{H}$, $\mathcal{K})$ and $L_{2}(\mathcal{H},\mathcal{K})$, we denote the linear spaces of bounded, compact, finite-rank and Hilbert–Schmidt operators between Hilbert spaces $\mathcal{H}$ and $\mathcal{K}$. The Hilbert–Schmidt norm of a linear operator between finite-dimensional spaces is equal to the Frobenius norm of any matrix representation of this operator in an orthonormal basis (ONB). By $\mathcal{B}_{00,r}(\mathcal{H},\mathcal{K})$ we denote the set of finite-rank operators, i.e. operators that are bounded, have finite-dimensional range, and have rank at most $r\in\mathbb{N}$. Further details of these classes of operators is given in Appendix A.
@@ 2 Optimal approximation for matrices
$\min{\left\{\lVert M-BXC\rVert_{F}:\ X\in\mathbb{C}^{p\times q},\ \operatorname{rank}\left({X}\right)\leq r\right\}},$ (4)
@@
A few remarks are in order. Firstly, in general not all solutions of (4) are of the form (3). Secondly, as $P_{C,R}=C^{\dagger}C$ and as in finite dimensions it holds that $P_{B,L}=BB^{\dagger}$, one may write (3) as $B^{\dagger}(BB^{\dagger}XC^{\dagger}C)_{r}C^{\dagger}$. Thirdly, also a necessary and sufficient condition for uniqueness of $\hat{X}$ is given in [20, Theorem 2.1]. Finally, if $(P_{B,L}MP_{C,R})_{r}$ is not uniquely specified, then the matrix $\hat{X}$ in (3) is not uniquely determined. Hence eq. 3 should be interpreted as a set of solutions. Furthermore, this lack of uniqueness extends also to $\lVert\hat{X}\rVert_{F}$: if there exists more than one solution of the form (3), then these solutions might not all have the same norm. In other words, whether $\hat{X}$ in (3) is of minimal norm or not in general depends on the choice of truncated SVD $(P_{B,L}MP_{C,R})_{r}$. This fact indicates that the claim above is not sufficiently precise.
@@
Hence $M=P_{B,L}MP_{C,R}$ and its rank-1 truncated SVDs are
@@
The two minimizing matrices of the form (3) then are
$\hat{X}_{a}\coloneqq\begin{pmatrix}1&0&0\\ 0&0&0\\ 0&0&0\\ \end{pmatrix},\quad\hat{X}_{b}\coloneqq\begin{pmatrix}0&0&0\\ 0&4&0\\ 0&0&0\\ \end{pmatrix}.$
@@
The preceding example suggests that [20, Theorem 2.1] can be made more precise, either by removing the minimum norm condition (5), or by replacing this condition by a weaker condition which all solutions $\hat{X}$ given by (3) satisfy. Such a weaker condition must thus be independent of the choice of truncated SVD $(P_{B,L}MP_{C,R})_{r}$. Furthermore, if $C$ is a linear operator mapping between infinite-dimensional spaces, then ${C}^{\dagger}$ will in general not be bounded (cf. Proposition A.5), and thus cannot be Hilbert–Schmidt. It can then happen that also solutions $\hat{X}$ of the form (3) are not Hilbert–Schmidt, in which case $\lVert\hat{X}\rVert_{L_{2}}$ may be ill-defined, and we need a minimality condition that does not involve the Frobenius norm in order to generalise this condition to the infinite-dimensional setting. For these two reasons we motivate an alternative condition, by continuing our Example 2.1.
@@
Like $B\hat{X}_{c}(\alpha)C$, also $BX(\alpha)C$ does not depend on $\alpha$. By the Eckart–Young Theorem (c.f. Proposition A.7), for $X(\alpha)$ to be a solution of (4), we need $BX(\alpha)C$ to be a rank-1 SVD of $M$. That is, we need either $x_{11}=1$ or $x_{22}=4$, and all other entries to be equal to zero.
@@
Let $e_{i}$ denote the $i$-th canonical basis vector of $\mathbb{C}^{3}$ and $V\coloneqq\operatorname{span}{\left(e_{1},e_{2}\right)}\subset\mathbb{C}^{3}$. Let $P_{V}$ the matrix representation of the projector onto $V$ in the canonical basis of $\mathbb{C}^{3}$. Then for $B,C$ defined in (6), $P_{B,R}=P_{V}=P_{C,L}$. We observe $P_{B,L}\hat{X}_{c}(\alpha)P_{C,R}=P_{V}\hat{X}_{c}(\alpha)P_{V}=\hat{X}_{c}(0)$ for any $\alpha$, and compute,
$\lVert\hat{X}_{c}(\alpha)\rVert_{F}^{2}=\lVert X_{c,11}\rVert_{F}^{2}+\sum_{i=1}^{5}\lvert\alpha_{i}\rvert^{2}\geq\lVert X_{c,11}\rVert_{F}^{2}=\lVert\hat{X}_{c}(0)\rVert_{F}^{2}.$ (7)
@@
that are the minimisers in (7) for some $c\in\{a,b\}$. Hence (8) provides an alternative for the minimal norm property (5). Thus, while $X_{a}(0)$ and $X_{b}(0)$ might not have the same norm, $X_{c}(0)$ has minimal norm among all $X_{c}(\alpha),\alpha\in\mathbb{C}^{5}$, for fixed $c\in\{a,b\}$, and $X_{a}(0),X_{b}(0)$ are in turn those solutions of (4) which satisfy (8).
To generalise the reasoning of Example 2.1 as to why (8) can be interpreted as a minimality condition, notice that (8) is equivalent to $(I-P_{B,R})\hat{X}P_{C,L}=0$ and $\hat{X}(I-P_{C,L})=0$, since any $X\in\mathbb{C}^{p\times q}$ can be decomposed as
@@
If $\hat{X}$ in (3) is unique, then $\lVert\hat{X}\rVert_{L_{2}}$ is indeed minimal among all minimisers $X$ of (4), in which case the statement of [20, Theorem 2.1] is still valid. Of course, this holds only when Frobenius norms can in fact be computed, which is no longer true in general for infinite-dimensional spaces. In the next section we introduce a statement similar to [20, Theorem 2.1] with the modified minimality property (8) in an infinite-dimensional setting.